# Concept Based Continuous Prompts for Interpretable Text Classification

### v1

@@ 6.1 Experimental Setups
Metrics and Baselines We investigate the interpretability of CD by comparing it to other explanation methods using two metrics: accuracy and concept correlation. Following Ju et al. (2023), we use accuracy to assess how well the concept prompt aligns with the continuous prompts. We adopt Discrete-1500 explanation methods as our baseline, with 1500 being the maximum number of vocabularies according to Ju et al. (2023). Inspired by Jin et al. (2020), we introduce a new metric called Concept Correlation, denoted as $\rho$, to evaluate concept-level explanations. Concretely, we compute the Pearson correlation between the coefficients learned by a bag-of-concepts model and the attribution scores provided by various explanation methods. For baselines, we choose three attribution methods to assign scores for each prompt. These scores are comparable to coefficients for explanation faithfulness. Specifically we choose TokenSHAP Goldshmidt and Horovicz (2024), IG Sundararajan et al. (2017), and grad Shrikumar et al. (2017).
@@
P-tuning CD Discrete P-tuning CD Discrete P-tuning CD P-tuning CD
  SST-2 4-shot 0.581 0.592 0.556 0.625 0.606 0.620 0.560 0.557 0.533 0.545
8-shot 0.691 0.615 0.618 0.579 0.674 0.642 0.636 0.596 0.605 0.603
16-shot 0.840 0.797 0.818 0.744 0.744 0.752 0.831 0.848 0.821 0.809
32-shot 0.847 0.850 0.859 0.831 0.812 0.782 0.839 0.852 0.832 0.856
@@
  IMDB 4-shot 0.542 0.577 0.556 0.563 0.538 0.569 0.547 0.556 0.529 0.529
8-shot 0.750 0.789 0.688 0.605 0.549 0.536 0.509 0.507 0.518 0.501
16-shot 0.834 0.876 0.852 0.848 0.772 0.759 0.506 0.507 0.504 0.527
32-shot 0.839 0.848 0.867 0.804 0.800 0.800 0.790 0.714 0.763 0.864
@@
  AGNews 4-shot 0.687 0.704 0.712 0.683 0.698 0.663 0.762 0.755 0.759 0.762
8-shot 0.742 0.692 0.689 0.712 0.711 0.741 0.790 0.790 0.797 0.796
16-shot 0.767 0.766 0.758 0.759 0.755 0.747 0.784 0.793 0.781 0.804
32-shot 0.823 0.814 0.828 0.828 0.837 0.805 0.836 0.837 0.831 0.831
@@
Table 2: Accuracy results of CD with baselines on the SST-2, IMDB and AGNews dataset. $N_{q}$ denotes the number of continuous prompt tokens. Discrete represents Discrete-1500. $\sigma^{2}$ denotes the average variance across five shot settings.
@@
CD TokenSHAP IG Grad CD TokenSHAP IG Grad
  SST-2 top-3 0.755 0.646 0.525 0.481 0.824 0.613 0.535 0.342
top-5 0.561 0.427 0.221 0.264 0.610 0.497 0.353 0.289
top-10 0.484 0.398 0.141 0.173 0.445 0.310 0.252 0.200
@@
  Dataset Class Input Top-3 Concepts
SST-2 positive I loved it! 1.Encouraging and Supportive. 2.Happy and Optimistic. 3.Constructive and Helpful.
negative extremely confusing. 1.Opposing or contradictory to desired outcomes. 2.Pessimistic outlook or attitude. 3.Discouraging or disheartening experiences.
  AGNews world euters china has said no date has been set for working level talks on the north korean nuclear crisis … 1.Breaking news updates on significant global events. 2.In-depth analysis of international conflicts and resolutions. 3.Coverage of diplomatic relations between countries.
@@
Table 4: Explanation examples on SST-2 and AGNews dataset with BERT-Large.

### v2

@@ 6.1 Experimental Setups
Metrics and Baselines We investigate the interpretability of CD by comparing it to other explanation methods using two metrics: accuracy and concept correlation. Following Ju et al. (2023), we use accuracy to assess how well the concept prompt aligns with the continuous prompts. We adopt Vocab-1500 explanation methods as our baseline, with 1500 being the maximum number of vocabularies according to Ju et al. (2023). Inspired by Jin et al. (2020), we introduce a new metric called Concept Correlation, denoted as $\rho$, to evaluate concept-level explanations. Concretely, we compute the Pearson correlation between the coefficients learned by a bag-of-concepts model and the attribution scores provided by various explanation methods. For baselines, we choose three attribution methods to assign scores for each prompt. These scores are comparable to coefficients for explanation faithfulness. Specifically we choose TokenSHAP Goldshmidt and Horovicz (2024), IG Sundararajan et al. (2017), and grad Shrikumar et al. (2017).
@@
Dataset   shot P-tuning CD Vocab P-tuning CD Vocab P-tuning CD P-tuning CD
  SST-2 4 0.581 0.592 0.556 0.625 0.606 0.620 0.560 0.557 0.533 0.545
8 0.691 0.615 0.618 0.579 0.674 0.642 0.636 0.596 0.605 0.603
16 0.840 0.797 0.818 0.744 0.744 0.752 0.831 0.848 0.821 0.809
32 0.847 0.850 0.859 0.831 0.812 0.782 0.839 0.852 0.832 0.856
@@
  IMDB 4 0.542 0.577 0.556 0.563 0.538 0.569 0.547 0.556 0.529 0.529
8 0.750 0.789 0.688 0.605 0.549 0.536 0.509 0.507 0.518 0.501
16 0.834 0.876 0.852 0.848 0.772 0.759 0.506 0.507 0.504 0.527
32 0.839 0.848 0.867 0.804 0.800 0.800 0.790 0.714 0.763 0.864
@@
  AGNews 4 0.687 0.704 0.712 0.683 0.698 0.663 0.762 0.755 0.759 0.762
8 0.742 0.692 0.689 0.712 0.711 0.741 0.790 0.790 0.797 0.796
16 0.767 0.766 0.758 0.759 0.755 0.747 0.784 0.793 0.781 0.804
32 0.823 0.814 0.828 0.828 0.837 0.805 0.836 0.837 0.831 0.831
@@
Table 2: Accuracy results of CD with baselines on the SST-2, IMDB and AGNews dataset. $N_{q}$ denotes the number of continuous prompt tokens. Vocab represents Vocab-1500. $\sigma^{2}$ denotes the average variance across five shot settings.
@@
Dataset top-k CD TokenSHAP IG Grad CD TokenSHAP IG Grad
  SST-2 3 0.755 0.646 0.525 0.481 0.824 0.613 0.535 0.342
5 0.561 0.427 0.221 0.264 0.610 0.497 0.353 0.289
10 0.484 0.398 0.141 0.173 0.445 0.310 0.252 0.200
@@
  Class Input Top-3 Concepts
world euters china has said no date has been set for working level talks on the north korean nuclear crisis … 1.Breaking news updates on significant global events. 2.In-depth analysis of international conflicts and resolutions. 3.Coverage of diplomatic relations between countries.
@@
Table 4: Explanation examples on AGNews dataset with BERT-Large. Top-3 concepts are ranked based on the coefficients in $\boldsymbol{Q}$.